# GME: Improving Universal Multimodal Retrieval by Multimodal LLMs

### v1

@@ 4.2 Fused-Modal Data Synthesis
To efficiently synthesize high-quality data while minimizing manual intervention, we adopt a strategy similar to Doc2Query (Gospodinov et al., 2023). However, our approach differs in that we aim to generate fuse-modal candidate-to-query relevance data instead of single-modality, text-based relevance pairs. This requires obtaining high-quality candidates that include both image and text content. We primarily extracted such data from Wikipedia paragraphs444WikiWeb2M. Additionally, to enhance the domain diversity of the candidate data, we employed a domain classification model555We use bart-large-mnli as the classifier. to perform fine-grained classification of Wikipedia data into categories such as animals and plants. We then uniformly sampled from these categories and retained data with classification confidence scores above 0.5. Ultimately, we obtained 313,284 candidate entries, each containing both text and image content.
Based on these prepared candidate data, as shown in Figure 4, the overall data synthesis pipeline can be divided into the following steps: Doc2Query Generation: The passage content from each candidate is input into an LLM666In the entire synthesis process, we utilize Qwen/Qwen2.5-72B-Instruct as our LLM. using a prompt to generate a natural query based on the passage. To ensure the quality of the generated queries, we built a vector index of all passage contents using a text vector retrieval model777We use Alibaba-NLP/gte-Qwen2-1.5B-instruct as the retriever.. Each generated query is then used to retrieve the corresponding passage from this collection. If the passage associated with the query is not within the top 20 retrieved items, the query is considered low quality due to low relevance and is discarded. In this step, we discarded 1.2% of the total generated queries. This process allows us to construct TIT training data. Entity Extraction and Query Rewrite: We aim for the synthesized queries to include both texts and images (i.e., ITIT type). To achieve this, we leverage entity extraction followed by image retrieval for the extracted entities and caption generation to supplement the image data on the query side. Specifically, for each generated query from the first step, we use an Entity Extraction approach with an LLM to extract entities and rewrite the original query into . For example, the query “Where is Iris pseudacorus native?” is transformed by the model to the rewritten query ”Where is the native habitat of this plant?” with the entity ”Iris pseudacorus” extracted. We then seek images that match this entity and combine them with the rewritten query to form the final fuse-modal query. Image Retrieval and Generation: We explore two methods for obtaining images. The first method uses the Google Image Search API888https://serpapi.com/google-images-api to retrieve images matching the entity terms, retaining the top five results. The second method involves generating images using a text-to-image model999https://huggingface.co/black-forest-labs/FLUX.1-dev. Specifically, we first use the LLM to generate a caption suitable for image generation based on the entity and the passage of the generated query, then input this caption into the text-to-image generation model to create the corresponding image. This approach allows us to quickly and efficiently obtain high-quality, diverse images. The synthesized results can also be assembled into IT IT retrieval type data. Data Filtering: To ensure the quality of the synthesized data, we performed filtering on the final dataset. We observed that images generated by the FLUX model had consistent quality, whereas images retrieved via the Google Image Search API often included noisy data. Therefore, for images obtained through the Google Image Search API, we used the CLIP model101010openai/clip-vit-large-patch14 to assess image-caption relevance. Images with a relevance score below 0.2 were filtered out.
@@
Through the aforementioned synthesis pipeline, we produced 1,135,000 high-quality fuse-modal training data entries (including TIT and ITIT types). After filtering, we retained 1,102,000 entries, resulting in a data loss rate of . The entire process consumed 600 A100 GPU hours. Detailed descriptions of all prompts used in the data synthesis pipeline and examples of the synthesized data will be provided in the Appendix §D.
@@
| VISTA (2024a) | 0.2B | 55.15 | 31.98 | 32.88 | 10.12 | 31.23 | 45.81 | 53.32 | 8.97 | 26.26 | 37.32 |
| CLIP-SF (2024) | 0.4B | 39.75 | 31.42 | 59.05 | 24.09 | 62.95 | 66.41 | 53.32 | 34.90 | 55.65 | 43.66 |
| One-Peace (2023) | 4B | 43.54 | 31.27 | 61.38 | 42.9 | 65.59 | 42.72 | 28.29 | 6.73 | 23.41 | 42.01 |
| DSE (2024) | 4.2B | 48.94 | 27.92 | 40.75 | 78.21 | 52.54 | 49.62 | 35.44 | 8.36 | 40.18 | 50.04 |
| E5-V (2024a) | 8.4B | 52.41 | 27.36 | 46.56 | 41.22 | 47.95 | 54.13 | 32.90 | 23.17 | 7.23 | 42.52 |
@@ 5.1 Settings
Building on the findings from Section 3, we trained our model using a diverse dataset of 8 million instances spanning various retrieval modalities. For single-modal retrieval tasks, we utilized datasets including MSMARCO (Nguyen et al., 2016), NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), TriviaQA (Joshi et al., 2017), SQuAD (Rajpurkar et al., 2016), FEVER (Thorne et al., 2018), and AllNLI for SimCSE (Gao et al., 2021), selecting a total of 1 million entries. From ImageNet (Deng et al., 2009), we extracted 1 million image-to-image training instances, designating images within the same class as positive samples and others as negative samples. For cross-modal retrieval tasks, we incorporated 2 million entries from the LAION (Schuhmann et al., 2022), mscoco (Lin et al., 2014), and Docmatix (Laurençon et al., 2024) datasets. Additionally, for fused-modal retrieval tasks, we included a total of 2 million instances: million synthesized by us, and the remaining from the M-BEIR (Wei et al., 2024) training data.
@@
We use Qwen2-VL (Wang et al., 2024b) model series as the backbone for our MLLM, conducting training on models with both 2 billion (2B) and 7 billion (7B) parameters. Our training utilizes Low-Rank Adaptation (LoRA) (Hu et al., 2022) with a rank of 8, a learning rate of 1e-4, and a temperature setting of 0.03. To manage the varying number of visual tokens required by Qwen2-VL for different image resolutions and maintain training efficiency, we limited the maximum number of visual tokens per image to 1,024.
For data with images, we set the maximum text length to 1,800 tokens, using a batch size of 128 for the 2B model and 32 for the 7B model. For text-only data, the maximum length was set to 512 tokens, with batch sizes of 512 for the 2B model and 128 for the 7B model. Each training sample included 8 negative examples. To conserve GPU memory, we employed gradient checkpointing (Chen et al., 2016) and trained the model using bfloat16 precision. All training was conducted on eight NVIDIA A100 GPUs, each with 80GB of memory.
@@
We compare our method against four types of retrieval systems: (1) Previous representative UMR models, for example, VISTA (Zhou et al., 2024a) for text encoder based, and E5-V (Jiang et al., 2024a) for MLLM based; (2) Powerful multimodal representation (embedding) models, i.e., One-Peace (Wang et al., 2023), which supports modalities beyond text and image and hence could also be tested on our UMRB; (3) Recent visual document retrieval models, namely DSE (Ma et al., 2024); and (4) the classic cross-modal pipeline, CLIP score-fusion, denoted as CLIP-SF, which provides top-tier cross-modal performance. We exclude comparisons with state-of-the-art text retrieval models as VISTA demonstrates comparable performance levels.
@@ 5.2 Main Results
Table 3 presents the evaluation results of the baseline systems alongside our proposed GME. Scores are averaged across each sub-task and categorized by retrieval modality type: single-modal, cross-modal, and fused-modal. Additionally, the overall micro-average score on the UMRB is in the last column. First, focusing on the average scores, our smaller model, i.e., GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA). The larger model, i.e., GME-Qwen2-VL-7B, further enhances this performance, demonstrating the effectiveness of our approach in handling UMR tasks.
Second, our models outperform smaller methods such as VISTA (million-level parameters) and One-Peace (4B parameters). The larger MLLM baseline, E5-V (8B parameters), performs well in text-dominated tasks (e.g., TT) but falls short in other areas. This indicates that training with multimodal data is crucial for achieving superior performance in UMR tasks. Our training data provides a stronger foundation for future advancements.
@@
Finally, we compare with the recent visual document retrieval model DSE, specialized for the TVD task within the Cross-Modal group, which has approximately 4B parameters. Our models are competitive with or exceed the performance of this task-specific baseline, demonstrating the feasibility and promise of integrating visual document retrieval into a unified retriever framework.
@@ 5.3 Analyses
In this subsection, we conduct several analyses to further understand our methods and models.
@@
Given our model’s impressive performance, we assesse the quality of its embeddings. Specifically, we investigate whether the embeddings are modality-universal meaning that embeddings representing the same semantic content across different modalities are closely clustered in the embedding space, or if they remain in separate sub-spaces tailored for each modality-specific task. To probe this question, we sample 1000 instances from the EVQA dataset and visualize their embeddings of different modalities by t-SNE, as shown in Figure 5. We can observe that the CLIP’s embeddings are distinctly separated by modality, whereas the embeddings from our model are intermingled and organized semantically. This demonstrates that our model effectively generates modality-universal representations, enhancing its applicability across various UMR tasks.
@@
Our approach is primarily data-centric, constructing a diverse training dataset of approximately million samples across various UMR settings (§5.1). Training on such a large-scale dataset demands significant computational resources and time, making arbitrary training decisions impractical. Therefore, we explored the training scaling law by examining how model performance evolves with increasing training steps. Due to the time-consuming nature of evaluating certain retrieval tasks, we assessed performance on our UMRB-Partial dataset for faster evaluation. Figure 6 illustrates the performance progression of our 2B and 7B models on UMRB-Partial during training. Both models exhibit linear performance improvements as training continues, suggesting that extended training could yield further benefits. However, due to time constraints, we halted current training. Future work will investigate longer training periods to enhance model performance further.
@@
We conduct an ablation study to investigate the effectiveness of different design choices of GME. We consider the following three aspects: (1) Fine-tuning strategy. Our final models are trained by LoRA with rank 8. We compare with other rank values and full fine-tuning. The results in the first group of Table 5 show that LoRA with rank 8 yields the best performance. (2) Training data organization. We compare models trained without hard negative mining or retrieval instructions. The second group of Table 5 demonstrates that both modifications led to performance declines, indicating that these two techniques are essential for effective model training. (3) Modeling techniques. Our final models are in the casual attention mode and use the EOS token state as the embedding, hence we compare the performance of the model trained with mean pooling and the bi-directional attention mechanism. The last group of Table 5 shows that these alternative settings negatively impact performance.
@@ 6 Conclusion
In this work, we target the universal multimodal retrieval (UMR) problem. We begin by systematically categorizing current UMR tasks, proposing a comprehensive classification framework. Based on this, we explore ways to further improve MLLM-based UMR models, suggesting the GME model. The GME models are trained using contrastive learning loss on a diverse set of multimodal data settings, while also extending support for visual retrieval. Additionally, to overcome limitations in existing UMR evaluation benchmarks, we compiled a new comprehensive benchmark (i.e., UMRB) by integrating multiple data sources. This benchmark effectively balances existing UMR tasks with the increasingly important text and visual document retrieval tasks, enabling a more thorough assessment of UMR model performance. We evaluate existing UMR models and our proposed GME model on UMRB, finding that our model achieves state-of-the-art performance. We also conducted various analyses to validate the effectiveness of our methods and enhance our understanding of them.
@@ Appendix A UMRB Details
A.1 Single-Modal Tasks
WebQA (Chang et al., 2022)
@@ A.1 Single-Modal Tasks
For these datasets, we use the processed versions from BEIR (Thakur et al., 2021).
A.2 Cross-Modal Tasks
VisualNews (Liu et al., 2021a)
@@ A.2 Cross-Modal Tasks
These datasets constitute the retrieval task of TVD. Their queries are standard questions, and the candidates are document screenshots. For these datasets, we used the processed versions from ViDoRe (Faysse et al., 2024b).
A.3 Fused-Modal Tasks
WebQA (Chang et al., 2022)
@@ A.3 Fused-Modal Tasks
ReMuQ (Luo et al., 2023)
The dataset is designed for a multimodal-query retriever, augmenting the WebQA questions by adding images to create new multimodal queries along with a large text corpus. For evaluation, we used 3,609 instances from this dataset.
OKVQA (Marino et al., 2019)

### v2

@@ 4.2 Fused-Modal Data Synthesis
To efficiently synthesize high-quality data while minimizing manual intervention, we adopt a strategy similar to Doc2Query [15]. However, our approach differs in that we aim to generate fuse-modal candidate-to-query relevance data instead of single-modality, text-based relevance pairs. This requires obtaining high-quality candidates that include both image and text content. We primarily extracted such data from Wikipedia paragraphs444github.com/google-research-datasets/wit/blob/main/wikiweb2m.md. Additionally, to enhance the domain diversity of the candidate data, we employed a domain classification model555hf.co/facebook/bart-large-mnlifacebook to perform fine-grained classification of Wikipedia data into categories such as animals and plants. We then uniformly sampled from these categories and retained data with classification confidence scores above 0.5. Ultimately, we obtained 313,284 candidate entries, each containing both text and image content.
Based on the prepared data, the overall synthesis pipeline (Figure 4) could be divided into the following steps:
Doc2Query Generation: The passage content from each candidate is input into an LLM666In the entire pipeline, we utilize Qwen2.5-72B-Instruct as our LLM. using a prompt to generate a natural query. To ensure the quality of the generated queries, we built a vector index of all passage contents using a text vector retrieval model777hf.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct. Each generated query is then used to retrieve the corresponding passage from this collection. If the passage associated with the query is not within the top 20 retrieved items, the query is considered low quality due to low relevance and is discarded. In this step, we discarded 1.2% of the total generated queries. This process allows us to construct TIT training data.
Entity Extraction and Query Rewrite: We aim for the synthesized queries to include both texts and images (i.e., ITIT type). To achieve this, we leverage entity extraction followed by image retrieval for the extracted entities and caption generation to supplement the image data on the query side. Specifically, for each generated query from the first step, we prompt the LLM to extract entities from it with the text passage as reference, and then rewrite the original query into . For example, the query “Where is Iris pseudacorus native?” is transformed by the model to the rewritten query ”Where is the native habitat of this plant?” with the entity ”Iris pseudacorus” extracted. We then seek images that match this entity and combine them with the rewritten query to form the final fuse-modal query.
Image Retrieval and Generation: We explore two methods for obtaining images. The first method uses the Google Image Search API888https://serpapi.com/google-images-api to retrieve images matching the entity terms, retaining the top five results. The second method involves generating images using a text-to-image model999 https://hf.co/black-forest-labs/FLUX.1-dev. Specifically, we first use the LLM to generate a caption suitable for image generation based on the entity and the passage of the generated query, then input this caption into the text-to-image generation model to create the corresponding image. This approach allows us to quickly and efficiently obtain high-quality, diverse images. The synthesized results can also be assembled into IT IT retrieval type data.
Data Filtering: To ensure the quality of the synthesized data, we perform filtering [9] on the final dataset. We observe that images generated by the FLUX model have consistent quality, whereas images retrieved via the Google Image Search API often include noisy data. Therefore, for images obtained through the Google Image Search API, we use the CLIP model101010https://hf.co/openai/clip-vit-large-patch14 to assess image-caption relevance. Images with a relevance score below were filtered out.
@@
Through the synthesis pipeline, we produce 1,135,000 high-quality fuse-modal training data entries (including TIT and ITIT types). After filtering, we retain 1,102,000 entries, resulting in a data loss rate of . The entire process consumed 600 A100 GPU hours. Detailed descriptions of all prompts used in the data synthesis pipeline and examples of the synthesized data are provided in the Appendix §10.
@@
| VISTA [74] | 0.2B | 55.15 | 31.98 | 32.88 | 10.12 | 31.23 | 45.81 | 53.32 | 8.97 | 26.26 | 37.32 |
| CLIP-SF [66] | 0.4B | 39.75 | 31.42 | 59.05 | 24.09 | 62.95 | 66.41 | 53.32 | 34.90 | 55.65 | 43.66 |
| One-Peace [64] | 4B | 43.54 | 31.27 | 61.38 | 42.9 | 65.59 | 42.72 | 28.29 | 6.73 | 23.41 | 42.01 |
| DSE [41] | 4.2B | 48.94 | 27.92 | 40.75 | 78.21 | 52.54 | 49.62 | 35.44 | 8.36 | 40.18 | 50.04 |
| E5-V [22] | 8.4B | 52.41 | 27.36 | 46.56 | 41.22 | 47.95 | 54.13 | 32.90 | 23.17 | 7.23 | 42.52 |
@@ 5.1 Settings
Building on the findings from §4.1, we train our model using a diverse dataset of 8 million instances spanning various retrieval modalities. For single-modal retrieval tasks, we utilize datasets including MSMARCO [47], NQ [26], HotpotQA [70], TriviaQA [24], SQuAD [52], FEVER [56], and AllNLI for SimCSE [14], selecting a total of 1 million entries. From ImageNet [10], we extract 1 million image-to-image training instances, designating images within the same class as positive samples and others as negative samples. For cross-modal retrieval tasks, we incorporate 2 million entries from the LAION [54], MSCOCO [32], and Docmatix [27] datasets. Additionally, for fused-modal retrieval tasks, we include a total of 2 million instances: million synthesized by us, and the remaining from the M-BEIR [66] training data.
@@
We use Qwen2-VL [65] model series as the backbone for our MLLM, conducting training on models with both 2 billion (2B) and 7 billion (7B) parameters. Our training utilizes Low-Rank Adaptation (LoRA) [19] with a rank of 8, a learning rate of 1e-4, and a temperature setting of 0.03. To manage the varying number of visual tokens required by Qwen2-VL for different image resolutions and maintain training efficiency, we limit the maximum number of visual tokens per image to 1,024.
For data with images, we set the maximum text length to 1,800 tokens, using a batch size of 128 for the 2B model and 32 for the 7B model. For text-only data, the maximum length was set to 512 tokens, with batch size of 512 for the 2B model and 128 for the 7B model. Each training sample included 8 negative examples. To conserve GPU memory, we employ gradient checkpointing [5] and train the model using bfloat16 precision. All training was conducted on eight NVIDIA A100 GPUs, each with 80GB of memory.
@@
We compare our method against four types of retrieval systems: (1) Previous representative UMR models, for example, VISTA [74] for text encoder based, and E5-V [22] for MLLM based; (2) Powerful multimodal representation (embedding) models, i.e. One-Peace [64], which supports modalities beyond text and image and hence could also be tested on our UMRB; (3) Recent visual document retrieval models, namely DSE [41]; and (4) the classic cross-modal pipeline, CLIP score-fusion, denoted as CLIP-SF, which provides top-tier cross-modal performance. We exclude comparisons with state-of-the-art text retrieval models as VISTA demonstrates comparable performance levels.
@@ 5.2 Main Results
Table 3 presents the evaluation results of the baseline systems alongside our proposed GME. Scores are averaged across each sub-task and categorized by retrieval modality type: single-modal, cross-modal, and fused-modal. Additionally, the overall micro-average score on the UMRB is in the last column. First, focusing on the average scores, our smaller model, i.e. GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA [74]). The larger model, i.e. GME-Qwen2-VL-7B, further enhances this performance, demonstrating the effectiveness of our approach in handling UMR tasks.
Second, our models outperform smaller methods such as VISTA (million-level parameters) and One-Peace (4B parameters). The larger MLLM baseline, E5-V [22] (8B parameters), performs well in text-dominated tasks (e.g., TT) but falls short in other areas. This indicates that training with multimodal data is crucial for achieving superior performance in UMR tasks. Our training data provides a stronger foundation for future advancements.
@@
Finally, we compare with the recent visual document retrieval model DSE [41], specialized for the TVD task within the Cross-Modal group, which has approximately 4B parameters. Our models are competitive with or exceed the performance of this task-specific baseline, demonstrating the feasibility and promise of integrating visual document retrieval into a unified retriever framework.
@@ 5.3 Analyses
Given our the impressive performance of our model, we assess the quality of its embeddings. Specifically, we investigate whether the embeddings are modality-universal meaning that embeddings representing the same semantic content across different modalities are closely clustered in the embedding space, or if they remain in separate sub-spaces tailored for each modality-specific task. To probe this question, we sample 1000 instances from the EVQA dataset and visualize their embeddings of different modalities by t-SNE, as shown in Figure 5. We also highlight two semantic close groups with yellow and pink labels, respectively. We can observe that the embeddings from CLIP are distinctly separated by modality, whereas the embeddings from our model are intermingled and organized semantically. Meanwhile, the points from the same semantic group are closely clustered. This demonstrates that our model effectively generates modality-universal representations, enhancing its applicability across various UMR tasks.
@@
Our approach is primarily data-centric, constructing a diverse training dataset of approximately million samples across various UMR settings (§5.1). Training on such a large-scale dataset demands significant computational resources and time. Therefore, we explored the training scaling law by examining how model performance evolves with increasing training steps. Due to the time-consuming nature of evaluating certain retrieval tasks, we assessed performance on our UMRB-Partial dataset for faster evaluation. Figure 6 illustrates the performance progression of our 2B and 7B models on UMRB-Partial during training. Both models exhibit linear performance improvements as training continues, suggesting that extended training could yield further benefits. However, due to time constraints, we halted current training. Future work will investigate longer training periods to enhance model performance further.
@@
We conduct an ablation study to investigate the effectiveness of different design choices of GME. We consider the following three aspects: (1) Fine-tuning strategy. Our final models are trained by LoRA with rank 8. We compare with other rank values and full fine-tuning. The results in the first group of Table 5 show that LoRA with rank 8 yields the best performance. (2) Training data organization. We compare models trained without hard negative mining. The second group of Table 5 demonstrates that the removal of hard negatives led to performance declines, indicating that it is essential for effective retrieval model training. (3) Retrieval instructions. We compare models trained without retrieval instructions. The third group shows that retrieval instructions are crucial for better UMR. (4) Modeling techniques. Our final models are in the casual attention mode and use the EOS token state as the embedding, hence we compare the performance of the model trained with mean pooling and the bi-directional attention mechanism. The last group of Table 5 shows that these alternative settings negatively impact performance.
@@ 6 Conclusion
In this work, we target the universal multimodal retrieval (UMR) problem. We begin by systematically categorizing current UMR tasks, proposing a comprehensive classification framework. Based on this, we explore ways to further improve MLLM-based UMR models, suggesting the GME model. The GME models are trained using contrastive learning loss on a diverse set of multimodal data settings, while also extending support for visual retrieval. Additionally, to overcome limitations in existing UMR evaluation benchmarks, we compiled a new comprehensive benchmark (i.e., UMRB) by integrating multiple data sources. This benchmark effectively balances existing UMR tasks with the increasingly important text and visual document retrieval tasks, enabling a more thorough assessment of UMR model performance. We evaluate existing UMR models and our proposed GME model on UMRB, finding that our model achieves state-of-the-art performance. We also conducted various analyses to validate the effectiveness of our methods and enhance our understanding of them. Our benchmark, models, and other materials are open-source at https://hf.co/Alibaba-NLP/gme-Qwen2-VL-7B-Instruct.
Acknowledgments
This work receives partial support from the Natural Science Foundation of China (under Grant 624B2048) and Research Grant Council of Hong Kong (PolyU/15209724).
@@ 7 UMRB Details
7.1 Single-Modal Tasks
WebQA [4]
@@ 7.1 Single-Modal Tasks
For these datasets, we use the processed versions from BEIR [55].
7.2 Cross-Modal Tasks
VisualNews [34]
@@ 7.2 Cross-Modal Tasks
These datasets constitute the retrieval task of TVD. Their queries are standard questions, and the candidates are document screenshots. For these datasets, we used the processed versions from ViDoRe [12].
7.3 Fused-Modal Tasks
WebQA [4]
@@ 7.3 Fused-Modal Tasks
ReMuQ [40]
The dataset is augmented from the WebQA questions by adding images to create new multimodal queries along with a large text corpus. For evaluation, we used 3,609 instances from this dataset.
OKVQA [43]